OpenStudy (kaylak12345):

what is the expression in factored form 2x2+16x+24

OpenStudy (kaylak12345):

@jabez177

OpenStudy (kaylak12345):

@mathmate

OpenStudy (mathmate):

Steps: first take out all visible common factors, in this case, 2. \(2x^2+16x+24=2(x^2+8x+12)\) We know that if we want to factor in the form (x+a)(x+b), then a*b=12, a+b=8. and a,b must both be factors of 12 (since a*b=12). I'll leave it to you to find a and b to finish the problem.

OpenStudy (kaylak12345):

2(x + 6)(x + 2) answer?

OpenStudy (kaylak12345):

@mathmate

OpenStudy (kaylak12345):

@mathmate

OpenStudy (vuriffy):

That is the correct answer.

OpenStudy (kaylak12345):

really yay

OpenStudy (vuriffy):

Yes, great job.

OpenStudy (mathmate):

@kaylak12345 Yes, correct. Well done! :)